# Hengwei Bian1,2,    Lingdong Kong1,3   Haozhe Xie4   Liang Pan1,†,‡   Yu Qiao1   Ziwei Liu4

### v2

@@ Abstract
Urban scene generation has been developing rapidly recently. However, existing methods primarily focus on generating static and single-frame scenes, overlooking the inherently dynamic nature of real-world driving environments. In this work, we introduce DynamicCity, a novel 4D occupancy generation framework capable of generating large-scale, high-quality dynamic 4D scenes with semantics. DynamicCity mainly consists of two key models. 1) A VAE model for learning HexPlane as the compact 4D representation. Instead of using naive averaging operations, DynamicCity employs a novel Projection Module to effectively compress 4D features into six 2D feature maps for HexPlane construction, which significantly enhances HexPlane fitting quality (up to $\mathbf{12.56}$ mIoU gain). Furthermore, we utilize an Expansion & Squeeze Strategy to reconstruct 3D feature volumes in parallel, which improves both network training efficiency and reconstruction accuracy than naively querying each 3D point (up to $\mathbf{7.05}$ mIoU gain, $\mathbf{2.06x}$ training speedup, and $\mathbf{70.84\%}$ memory reduction). 2) A DiT-based diffusion model for HexPlane generation. To make HexPlane feasible for DiT generation, a Padded Rollout Operation is proposed to reorganize all six feature planes of the HexPlane as a squared 2D feature map. In particular, various conditions could be introduced in the diffusion or sampling process, supporting versatile 4D generation applications, such as trajectory- and command-driven generation, inpainting, and layout-conditioned generation. Extensive experiments on the CarlaSC and Waymo datasets demonstrate that DynamicCity significantly outperforms existing state-of-the-art 4D occupancy generation methods across multiple metrics. The code and models have been released to facilitate future research.
Figure 1: Dynamic occupancy generation from DynamicCity. We introduce a new generation model that generates diverse 4D scenes of large spatial scales ($80\times 80\times 6.4~{}\text{meter}^{3}$) and long sequential modeling (up to $128$ frames), enabling a diverse set of downstream applications. For more detailed examples, kindly refer to our Project Page: https://dynamic-city.github.io.
@@ 1 Introduction
Urban scene generation has garnered growing attention recently, which could benefit various related applications, such as robotics and autonomous driving. Compared to its 3D object generation counterpart, generating urban scenes remains an under-explored field, with many new research challenges such as the presence of numerous moving objects, large-scale scenes, and long temporal sequences (Huang et al., 2021). For example, in autonomous driving scenarios, a scene typically comprises multiple objects from various categories, such as vehicles, pedestrians, and vegetation, captured over a long sequence (e.g., $200$ frames) spanning a large area (e.g., $80\times 80\times 6.4~{}\text{meters}^{3}$). Although in its early stage, 4D occupancy generation holds great potential to enhance the understanding of the 3D world, with wide-reaching and profound implications.
Due to the complexity of occupancy data, many efficient learning frameworks have been introduced for large-scale 3D scene generation. $\mathcal{X}^{3}$ (Ren et al., 2024b) utilizes a hierarchical voxel diffusion model to generate outdoor 3D scenes based on VDB data structure. PDD (Liu et al., 2023a) introduces a pyramid discrete diffusion model to progressively generate high-quality 3D scenes. SemCity (Lee et al., 2024) resolves outdoor scene generation by leveraging a triplane diffusion model. Despite achieving impressive occupancy generation, they primarily focus on generating static and single-frame 3D occupancy, and hence fail to effectively capture the dynamic nature of outdoor environments. Recently, a few works (Zheng et al., 2024b; Wang et al., 2024) have explored 4D scene generation. However, generating high-quality long-sequence 4D scenes is still a challenging and open problem (Nakashima & Kurazume, 2021; Nakashima et al., 2023).
In this work, we propose a novel 4D occupancy generation framework, DynamicCity, enabling generating large-scale, high-quality dynamic occupancy scenes, which mainly consists of two stages: 1) a VAE network for learning compact 4D representations, i.e., HexPlanes (Cao & Johnson, 2023; Fridovich-Keil et al., 2023); 2) a HexPlane Generation model based on DiT (Peebles & Xie, 2023).
@@
  • In the VAE architecture, DynamicCity employs a novel Projection Module to benefit in encoding 4D scenes into compact HexPlanes, significantly improving HexPlane fitting quality. Following, an Expansion &\&& Squeeze Strategy is introduced to decode the HexPlanes for reconstruction, which improves both fitting efficiency and accuracy.
@@ 2 Related Work
3D Object Generation has been a central focus in machine learning, with diffusion models playing a significant role in generating realistic 3D structures. Many techniques utilize 2D diffusion mechanisms to synthesize 3D outputs, covering tasks like text-to-3D object generation (Ma et al., 2024), image-to-3D transformations (Wu et al., 2024a), and 3D editing (Rojas et al., 2024). Meanwhile, recent methods bypass the reliance on 2D intermediaries by generating 3D outputs directly in three-dimensional space, utilizing explicit (Alliegro et al., 2023), implicit (Liu et al., 2023b), triplane (Wu et al., 2024b), and latent representations (Ren et al., 2024b). Although these methods demonstrate impressive 3D object generation, they primarily focus on small-scale, isolated objects rather than large-scale, scene-level generation (Hong et al., 2024; Xie et al., 2025b). This limitation underscores the need for methods capable of generating complete 3D scenes with complex spatial relationships.
Urban Scene Generation extends the scope to larger, more complex environments. Earlier works used VQ-VAE (Zyrianov et al., 2022) and GAN-based models (Caccia et al., 2019; Nakashima et al., 2023) to generate LiDAR scenes. However, recent advancements have shifted towards diffusion models (Xiong et al., 2023; Ran et al., 2024; Nakashima & Kurazume, 2024; Zyrianov et al., 2022; Hu et al., 2024; Nunes et al., 2024), which better handle the complexities of expansive outdoor scenes. For example, (Lee et al., 2024) utilize voxel grids to represent large-scale scenes but often face challenges with empty spaces like skies and fields. While some recent works incorporate temporal dynamics to extend single-frame generation to sequences (Zheng et al., 2024b; Wang et al., 2024), they often lack the ability to fully capture the dynamic nature of 4D environments. Thus, these methods typically remain limited to short temporal horizons or struggle with realistic dynamic object modeling, highlighting the gap in generating high-fidelity 4D scenes.
4D Generation represents a leap forward, aiming to capture the temporal evolution of scenes. Prior works often leverage video diffusion models (Singer et al., 2022; Blattmann et al., 2023) to generate dynamic sequences (Singer et al., 2023), with some extending to multi-view (Shi et al., 2023; Xie et al., 2025a) and single-image settings (Rombach et al., 2022) to enhance 3D consistency. In the context of video-conditional generation, approaches such as (Jiang et al., 2023; Ren et al., 2023; 2024a) incorporate image priors for guiding generation processes. While these methods capture certain dynamic aspects, they lack the ability to generate long-term, high-resolution 4D scenes with versatile applications. Our method, DynamicCity, fills this gap by introducing a novel 4D generation framework that efficiently captures large-scale dynamic environments, supports diverse generation tasks (e.g., trajectory-guided (Bahmani et al., 2024), command-driven generation), and offers substantial improvements in scene fidelity and temporal modeling.
@@ 3 Preliminaries
HexPlane (Cao & Johnson, 2023; Fridovich-Keil et al., 2023) is an explicit and structured representation designed for efficient modeling of dynamic 3D scenes, leveraging feature planes to encode spacetime data. A dynamic 3D scene is represented as six 2D feature planes, each aligned with one of the major planes in the 4D spacetime grid. These planes are represented as $\mathcal{H}=[\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},\mathcal{P}_{tz}]$, comprising a Spatial TriPlane (Chan et al., 2022) with $\mathcal{P}_{xy}$, $\mathcal{P}_{xz}$, and $\mathcal{P}_{yz}$, and a Spatial-Time TriPlane with $\mathcal{P}_{tx}$, $\mathcal{P}_{ty}$, and $\mathcal{P}_{tz}$. To query the HexPlane at a point $\mathbf{p}=(t,x,y,z)$, features are extracted from the corresponding coordinates on each of the six planes and fused into a comprehensive representation. This fused feature vector is then passed through a lightweight network to predict scene attributes for $\mathbf{p}$.
Diffusion Transformers (DiT) (Peebles & Xie, 2023) are diffusion-based generative models using transformers to gradually convert Gaussian noise into data samples through denoising steps. The forward diffusion adds Gaussian noise over time, with a noised sample at step $t$ given by $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\alpha_{t}$ controls the noise schedule. The reverse diffusion, using a neural network $\epsilon_{\theta}$, aims to denoise $\mathbf{x}_{t}$ to recover $\mathbf{x}_{0}$, expressed as: $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t))$. New samples are generated by repeating this reverse process.
@@ 4 Our Approach
DynamicCity strives to generate dynamic 3D occupancy with semantic information, which mainly consists of a VAE for 4D occupancy encoding using HexPlane (Cao & Johnson, 2023; Fridovich-Keil et al., 2023) (Sec. 4.1), and a DiT for HexPlane generation (Sec. 4.2). Given a 4D scene, i.e., a dynamic 3D occupancy sequence $\mathbf{Q}\in\mathbb{R}^{T\times X\times Y\times Z\times C}$, where $T$, $X$, $Y$, $Z$, and $C$ denote the sequence length, height, width, depth, and channel size, respectively, the VAE first aims to encode an efficient 4D representation, HexPlane $\mathcal{H}=[\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},\mathcal{P}_{tz}]$, which is then decoded for reconstructing 4D scenes with semantics. After obtaining HexPlane embeddings, DynamicCity leverages a DiT-based framework for 4D occupancy generation. Diverse conditions could be introduced into the generation process, facilitating a range of downstream applications (Sec. 4.3). The overview of the proposed DynamicCity pipeline is illustrated in Fig. 2.
@@ 4.1 VAE for 4D Occupancy
To encode and compress $\mathcal{X}_{txyz}$ into compact 2D feature maps of $\mathcal{H}$, we propose a novel Projection Module with multiple projection networks $\bm{h}(\cdot)$. To project a high-dimensional feature input $\mathcal{X}_{\text{in}}\in\mathbb{R}^{{D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}}\times{D_{\text{r}}^{1}\times D_{\text{r}}^{2}\times\cdots\times D_{\text{r}}^{m}}\times C}$ as a lower-dimensional feature output $\mathcal{X}_{\text{out}}\in\mathbb{R}^{{D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}}\times C}$, the projection network $\bm{h}_{S_{\text{r}}}(\cdot)$ first reshapes $\mathcal{X}_{in}$ into a 3-dimensional feature $\mathcal{X}^{\prime}_{S_{\text{k}}S_{\text{r}}}\in\mathbb{R}^{S_{\text{k}}\times S_{\text{r}}\times C}$ by grouping the dimensions into the two new dimensions, i.e., $S_{\text{k}}$ the dimension that will be kept, and $S_{\text{r}}$ the dimension that will be reduced, where $S_{\text{k}}=D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}$, and $S_{\text{r}}=D_{\text{r}}^{1}\times D_{\text{r}}^{2}\times\cdots\times D_{\text{r}}^{m}$. Afterward, $\bm{h}_{S_{\text{r}}}(\cdot)$ utilizes a transformer-based operation to project the reshaped feature $\mathcal{X}^{\prime}_{S_{\text{k}}S_{\text{r}}}$ to $\mathcal{X}^{\prime\prime}_{S_{\text{k}}}\in\mathbb{R}^{S_{\text{k}}\times C}$, which is then reshaped to the expected lower-dimensional feature output $\mathcal{X}_{\text{out}}$. Formally, the projection network is formulated as:
$\mathcal{X}_{\text{out}}^{\{D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}\}\times C}=\bm{h}_{S_{\text{r}}}(\mathcal{X}_{\text{in}}^{\{D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}\}\times\{D_{\text{r}}^{1}\times D_{\text{r}}^{2}\times\cdots\times D_{\text{r}}^{m}\}\times C})~{},$ (1)
@@
To construct the spatial feature planes $\mathcal{P}_{xy}$, $\mathcal{P}_{xz}$, and $\mathcal{P}_{yz}$, the Projection Module first generates the XYZ Feature Volume $\mathcal{X}_{xyz}=\bm{h}_{t}(\mathcal{X}_{txyz})$. Rather than directly access the heavy feature volume sequence $\mathcal{X}_{txyz}$, $\bm{h}_{z}(\cdot)$, $\bm{h}_{y}(\cdot)$, and $\bm{h}_{x}(\cdot)$ are applied to $\mathcal{X}_{xyz}$ for reducing the spatial dimensions of $\mathcal{X}_{xyz}$ along the $z$-axis, $y$-axis, and $x$-axis, respectively. The temporal feature planes $\mathcal{P}_{tx},\mathcal{P}_{ty},$ and $\mathcal{P}_{tz}$ are directly obtained from $\mathcal{X}_{txyz}$ by simultaneously removing two spatial dimensions with $\bm{h}_{zy}(\cdot),\bm{h}_{xz}(\cdot)$, and $\bm{h}_{xy}(\cdot)$, respectively. Consequently, we could construct the HexPlane $\mathcal{H}$ based on the encoded six feature planes, including $\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},$ and $\mathcal{P}_{tz}$.
Decoding HexPlane. Based on the HexPlane $\mathcal{H}=[\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},\mathcal{P}_{tz}]$, we employ an Expansion & Squeeze Strategy (ESS), which could efficiently recover the feature volume sequence by decoding the feature planes in parallel for 4D occupancy reconstruction. ESS first duplicates and expands each feature plane $\mathcal{P}$ to match the shape of $\mathcal{X}_{txyz}$, resulting in the list of six feature volume sequences: $\{\mathcal{X}_{txyz}^{{P}_{xy}},\mathcal{X}_{txyz}^{{P}_{xz}},\mathcal{X}_{txyz}^{{P}_{yz}},\mathcal{X}_{txyz}^{{P}_{tx}},\mathcal{X}_{txyz}^{{P}_{ty}},\mathcal{X}_{txyz}^{{P}_{tz}}\}$ Afterward, ESS squeezes the corresponding six expanded feature volumes with Hadamard Product:
$\mathcal{X}^{\prime}_{txyz}=\prod_{\text{Hadamard}}{\{\mathcal{X}_{txyz}^{{P}_{xy}},\mathcal{X}_{txyz}^{{P}_{xz}},\mathcal{X}_{txyz}^{{P}_{yz}},\mathcal{X}_{txyz}^{{P}_{tx}},\mathcal{X}_{txyz}^{{P}_{ty}},\mathcal{X}_{txyz}^{{P}_{tz}}\}}~{}.$ (2)
@@
$\mathbf{Q}^{\prime}=g_{\phi}(\texttt{Concat}(\mathcal{X}^{\prime}_{txyz},\texttt{PE}(\texttt{Pos}(\mathcal{X}^{\prime}_{txyz}))))~{},$ (3)
@@
Optimization. The VAE is trained with a combined loss $\mathcal{L}_{\text{VAE}}$, including a cross-entropy loss, a Lovász-softmax loss (Berman et al., 2018), and a Kullback-Leibler (KL) divergence loss:
$\mathcal{L}_{\text{VAE}}=\mathcal{L}_{\text{CE}}(\mathbf{Q},\mathbf{Q}^{\prime})+\alpha\mathcal{L}_{\text{Lov}}(\mathbf{Q},\mathbf{Q}^{\prime})+\beta\mathcal{L}_{\text{KL}}(\mathcal{H},\mathcal{N}(\mathbf{0},\mathbf{I}))~{},$ (4)

### v3

# ] [Uncaptioned image] WorldBench Team

@@ Abstract
Urban scene generation has been developing rapidly recently. However, existing methods primarily focus on generating static and single-frame scenes, overlooking the inherently dynamic nature of real-world driving environments. In this work, we introduce DynamicCity, a novel 4D occupancy generation framework capable of generating large-scale, high-quality dynamic 4D scenes with semantics. DynamicCity mainly consists of two key models. (1) A VAE model for learning HexPlane as the compact 4D representation. Instead of using naive averaging operations, DynamicCity employs a novel Projection Module to effectively compress 4D features into six 2D feature maps for HexPlane construction, which significantly enhances HexPlane fitting quality (up to $\mathbf{12.56}$ mIoU gain). Furthermore, we utilize an Expansion & Squeeze Strategy to reconstruct 3D feature volumes in parallel, which improves both network training efficiency and reconstruction accuracy than naively querying each 3D point (up to $\mathbf{7.05}$ mIoU gain, $\mathbf{2.06x}$ training speedup, and $\mathbf{70.84\%}$ memory reduction). (2) A DiT-based diffusion model for HexPlane generation. To make HexPlane feasible for DiT generation, a Padded Rollout Operation is proposed to reorganize all six feature planes of the HexPlane as a squared 2D feature map. In particular, various conditions could be introduced in the diffusion or sampling process, supporting versatile 4D generation applications, such as trajectory- and command-driven generation, inpainting, and layout-conditioned generation. Extensive experiments on the CarlaSC and Waymo datasets demonstrate that DynamicCity significantly outperforms existing state-of-the-art 4D occupancy generation methods across multiple metrics. The code and models have been released to facilitate future research.
[…]
[ [Uncaptioned image] Project Page]https://dynamic-city.github.io
[…]
Figure 1: Dynamic occupancy generation from DynamicCity. We introduce a new 4D generation model that generates diverse 4D scenes of large spatial scales ($80\times 80\times 6.4\penalty 10000\ \text{meter}^{3}$) and long sequential modeling (up to $128$ frames), enabling a diverse set of downstream applications. For more detailed examples, kindly refer to our Project Page.
@@ 1 Introduction
Urban scene generation has garnered growing attention recently, which could benefit various related applications, such as robotics and autonomous driving. Compared to its 3D object generation counterpart, generating urban scenes remains an under-explored field, with many new research challenges such as the presence of numerous moving objects, large-scale scenes, and long temporal sequences [15]. For example, in autonomous driving scenarios, a scene typically comprises multiple objects from various categories, such as vehicles, pedestrians, and vegetation, captured over a long sequence (e.g., $200$ frames) spanning a large area (e.g., $80\times 80\times 6.4\penalty 10000\ \text{meters}^{3}$). Although in its early stages, 4D occupancy generation holds great potential to enhance the understanding of the 3D world, with wide-reaching and profound implications.
Due to the complexity of occupancy data, many efficient learning frameworks have been introduced for large-scale 3D scene generation. $\mathcal{X}^{3}$ [30] utilizes a hierarchical voxel diffusion model to generate outdoor 3D scenes based on VDB data structure. PDD [18] introduces a pyramid discrete diffusion model to progressively generate high-quality 3D scenes. SemCity [17] resolves outdoor scene generation by leveraging a triplane diffusion model. Despite achieving impressive occupancy generation, they primarily focus on generating static and single-frame 3D occupancy, and hence fail to effectively capture the dynamic nature of outdoor environments. Recently, a few works [49, 41] have explored 4D scene generation. However, generating high-quality long-sequence 4D scenes is still a challenging and open problem [21, 23].
In this work, we propose a novel 4D occupancy generation framework, DynamicCity, enabling the generation of large-scale, high-quality dynamic occupancy scenes, which mainly consists of two stages:
  • A VAE network for learning compact 4D representations, i.e., HexPlanes [7, 11].
  • A HexPlane Generation model based on DiT [26].
@@
  • In the VAE architecture, DynamicCity employs a novel Projection Module to benefit in encoding 4D scenes into compact HexPlanes, significantly improving HexPlane fitting quality. Following, an Expansion &\& Squeeze Strategy is introduced to decode the HexPlanes for reconstruction, which improves both fitting efficiency and accuracy.
@@ 2 Related Work
3D Object Generation has been a central focus in machine learning, with diffusion models playing a significant role in generating realistic 3D structures. Many techniques utilize 2D diffusion mechanisms to synthesize 3D outputs, covering tasks like text-to-3D object generation [20], image-to-3D transformations [43], and 3D editing [31]. Meanwhile, recent methods bypass the reliance on 2D intermediaries by generating 3D outputs directly in three-dimensional space, utilizing explicit [1], implicit [19], triplane [44], and latent representations [30]. Although these methods demonstrate impressive 3D object generation, they primarily focus on small-scale, isolated objects rather than large-scale, scene-level generation [13, 46]. This limitation underscores the need for methods capable of generating complete 3D scenes with complex spatial relationships.
Urban Scene Generation extends the scope to larger, more complex environments. Earlier works used VQ-VAE [50] and GAN-based models [5, 23] to generate LiDAR scenes. However, recent advancements have shifted towards diffusion models [47, 27, 22, 50, 14, 24], which better handle the complexities of expansive outdoor scenes. For example, [17] utilize voxel grids to represent large-scale scenes but often face challenges with empty spaces like skies and fields. While some recent works incorporate temporal dynamics to extend single-frame generation to sequences [49, 41], they often lack the ability to fully capture the dynamic nature of 4D environments. Thus, these methods typically remain limited to short temporal horizons or struggle with realistic dynamic object modeling, highlighting the gap in generating high-fidelity 4D scenes.
4D Generation represents a leap forward, aiming to capture the temporal evolution of scenes. Prior works often leverage video diffusion models [35, 4] to generate dynamic sequences [36], with some extending to multi-view [33, 45] and single-image settings [32] to enhance 3D consistency. In the context of video-conditional generation, approaches such as [16, 28, 29] incorporate image priors for guiding generation processes. While these methods capture certain dynamic aspects, they lack the ability to generate long-term, high-resolution 4D scenes with versatile applications. Our method, DynamicCity, fills this gap by introducing a novel 4D generation framework that efficiently captures large-scale dynamic environments, supports diverse generation tasks (e.g., trajectory-guided [2], command-driven generation), and offers substantial improvements in scene fidelity and temporal modeling.
@@ 3 Preliminaries
HexPlane [7, 11] is an explicit and structured representation designed for efficient modeling of dynamic 3D scenes, leveraging feature planes to encode spacetime data. A dynamic 3D scene is represented as six 2D feature planes, each aligned with one of the major planes in the 4D spacetime grid. These planes are represented as $\mathcal{H}=[\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},\mathcal{P}_{tz}]$, comprising a Spatial TriPlane [8] with $\mathcal{P}_{xy}$, $\mathcal{P}_{xz}$, and $\mathcal{P}_{yz}$, and a Spatial-Time TriPlane with $\mathcal{P}_{tx}$, $\mathcal{P}_{ty}$, and $\mathcal{P}_{tz}$. To query the HexPlane at a point $\mathbf{p}=(t,x,y,z)$, features are extracted from the corresponding coordinates on each of the six planes and fused into a comprehensive representation. This fused feature vector is then passed through a lightweight network to predict scene attributes for $\mathbf{p}$.
Diffusion Transformers (DiT) [26] are diffusion-based generative models using transformers to gradually convert Gaussian noise into data samples through denoising steps. The forward diffusion adds Gaussian noise over time, with a noised sample at step $t$ given by $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{1-\alpha_{t}}\mathbf{\epsilon},\quad\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\alpha_{t}$ controls the noise schedule. The reverse diffusion, using a neural network $\epsilon_{\theta}$, aims to denoise $\mathbf{x}_{t}$ to recover $\mathbf{x}_{0}$, expressed as: $\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(\mathbf{x}_{t},t))$. New samples are generated by repeating this reverse process.
@@ 4 Our Approach
DynamicCity strives to generate dynamic 3D occupancy with semantic information, which mainly consists of a VAE for 4D occupancy encoding using HexPlane [7, 11] (Sec. 4.1), and a DiT for HexPlane generation (Sec. 4.2). Given a 4D scene, i.e., a dynamic 3D occupancy sequence $\mathbf{Q}\in\mathbb{R}^{T\times X\times Y\times Z\times C}$, where $T$, $X$, $Y$, $Z$, and $C$ denote the sequence length, height, width, depth, and channel size, respectively, the VAE first aims to encode an efficient 4D representation, HexPlane $\mathcal{H}=[\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},\mathcal{P}_{tz}]$, which is then decoded for reconstructing 4D scenes with semantics.
After obtaining HexPlane embeddings, DynamicCity leverages a DiT-based framework for 4D occupancy generation. Diverse conditions could be introduced into the generation process, facilitating a range of downstream applications (Sec. 4.3). The overview of the proposed DynamicCity pipeline is illustrated in Fig. 2.
@@ 4.1 VAE for 4D Occupancy
To encode and compress $\mathcal{X}_{txyz}$ into compact 2D feature maps of $\mathcal{H}$, we propose a novel Projection Module with multiple projection networks $\bm{h}(\cdot)$.
To project a high-dimensional feature input $\mathcal{X}_{\text{in}}\in\mathbb{R}^{{D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}}\times{D_{\text{r}}^{1}\times D_{\text{r}}^{2}\times\cdots\times D_{\text{r}}^{m}}\times C}$ as a lower-dimensional feature output $\mathcal{X}_{\text{out}}\in\mathbb{R}^{{D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}}\times C}$, the projection network $\bm{h}_{S_{\text{r}}}(\cdot)$ first reshapes $\mathcal{X}_{in}$ into a 3-dimensional feature $\mathcal{X}^{\prime}_{S_{\text{k}}S_{\text{r}}}\in\mathbb{R}^{S_{\text{k}}\times S_{\text{r}}\times C}$ by grouping the dimensions into the two new dimensions, i.e., $S_{\text{k}}$ the dimension that will be kept, and $S_{\text{r}}$ the dimension that will be reduced, where $S_{\text{k}}=D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}$, and $S_{\text{r}}=D_{\text{r}}^{1}\times D_{\text{r}}^{2}\times\cdots\times D_{\text{r}}^{m}$. Afterward, $\bm{h}_{S_{\text{r}}}(\cdot)$ utilizes a transformer-based operation to project the reshaped feature $\mathcal{X}^{\prime}_{S_{\text{k}}S_{\text{r}}}$ to $\mathcal{X}^{\prime\prime}_{S_{\text{k}}}\in\mathbb{R}^{S_{\text{k}}\times C}$, which is then reshaped to the expected lower-dimensional feature output $\mathcal{X}_{\text{out}}$. Formally, the projection network is formulated as:
$\mathcal{X}_{\text{out}}^{\{D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}\}\times C}=\bm{h}_{S_{\text{r}}}(\mathcal{X}_{\text{in}}^{\{D_{\text{k}}^{1}\times D_{\text{k}}^{2}\times\cdots\times D_{\text{k}}^{n}\}\times\{D_{\text{r}}^{1}\times D_{\text{r}}^{2}\times\cdots\times D_{\text{r}}^{m}\}\times C})\penalty 10000\ ,$ (1)
@@
To construct the spatial feature planes $\mathcal{P}_{xy}$, $\mathcal{P}_{xz}$, and $\mathcal{P}_{yz}$, the Projection Module first generates the XYZ Feature Volume $\mathcal{X}_{xyz}=\bm{h}_{t}(\mathcal{X}_{txyz})$. Rather than directly access the heavy feature volume sequence $\mathcal{X}_{txyz}$, $\bm{h}_{z}(\cdot)$, $\bm{h}_{y}(\cdot)$, and $\bm{h}_{x}(\cdot)$ are applied to $\mathcal{X}_{xyz}$ for reducing the spatial dimensions of $\mathcal{X}_{xyz}$ along the z-axis, y-axis, and x-axis, respectively. The temporal feature planes $\mathcal{P}_{tx},\mathcal{P}_{ty},$ and $\mathcal{P}_{tz}$ are directly obtained from $\mathcal{X}_{txyz}$ by simultaneously removing two spatial dimensions with $\bm{h}_{zy}(\cdot),\bm{h}_{xz}(\cdot)$, and $\bm{h}_{xy}(\cdot)$, respectively.
Consequently, we could construct the HexPlane $\mathcal{H}$ based on the encoded six feature planes, including $\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},$ and $\mathcal{P}_{tz}$.
Decoding HexPlane. Based on the HexPlane $\mathcal{H}=[\mathcal{P}_{xy},\mathcal{P}_{xz},\mathcal{P}_{yz},\mathcal{P}_{tx},\mathcal{P}_{ty},\mathcal{P}_{tz}]$, we employ an Expansion & Squeeze Strategy (ESS), which could efficiently recover the feature volume sequence by decoding the feature planes in parallel for 4D occupancy reconstruction.
ESS first duplicates and expands each feature plane $\mathcal{P}$ to match the shape of $\mathcal{X}_{txyz}$, resulting in the list of six feature volume sequences: $\{\mathcal{X}_{txyz}^{{P}_{xy}},\mathcal{X}_{txyz}^{{P}_{xz}},\mathcal{X}_{txyz}^{{P}_{yz}},\mathcal{X}_{txyz}^{{P}_{tx}},\mathcal{X}_{txyz}^{{P}_{ty}},\mathcal{X}_{txyz}^{{P}_{tz}}\}$ . Afterward, ESS squeezes the corresponding six expanded feature volumes with the Hadamard Product:
$\mathcal{X}^{\prime}_{txyz}=\prod_{\text{Hadamard}}{\{\mathcal{X}_{txyz}^{{P}_{xy}},\mathcal{X}_{txyz}^{{P}_{xz}},\mathcal{X}_{txyz}^{{P}_{yz}},\mathcal{X}_{txyz}^{{P}_{tx}},\mathcal{X}_{txyz}^{{P}_{ty}},\mathcal{X}_{txyz}^{{P}_{tz}}\}}\penalty 10000\ .$ (2)
@@
$\mathbf{Q}^{\prime}=g_{\phi}(\texttt{Concat}(\mathcal{X}^{\prime}_{txyz},\texttt{PE}(\texttt{Pos}(\mathcal{X}^{\prime}_{txyz}))))\penalty 10000\ ,$ (3)
@@
Optimization. The VAE is trained with a combined loss $\mathcal{L}_{\text{VAE}}$, including a cross-entropy loss, a Lovász-softmax loss [3], and a Kullback-Leibler (KL) divergence loss:
$\mathcal{L}_{\text{VAE}}=\mathcal{L}_{\text{CE}}(\mathbf{Q},\mathbf{Q}^{\prime})+\alpha\mathcal{L}_{\text{Lov}}(\mathbf{Q},\mathbf{Q}^{\prime})+\beta\mathcal{L}_{\text{KL}}(\mathcal{H},\mathcal{N}(\mathbf{0},\mathbf{I}))\penalty 10000\ ,$ (4)